A gas cylinder is filled with argon at a pressure of 177atm and 25•C. What is the gas pressure when the temperature of the cylinder and its contents are heated to 195•C by exposure to fire?
1
Expert's answer
2019-12-03T08:24:37-0500

pV=nRT

p1/p2=T1/T2

p2=p1×T2/T1

p2=278 atm